The area of a rectangle is 48cm^2 ,the length is halved and the width is tripled. What is the new area in cm^2
Answer by ankor@dixie-net.com(22740)

The area of a rectangle is 48cm^2,
L*W = 48
the length is halved and the width is tripled.
.5L * 3w = 1.5Lw
What is the new area in cm^2
Replace Lw with 48
1.5 * 48 = 72 sq/cm is the new area